OpenStudy (anonymous):

Which of the following expressions demonstrates the distributive property? A.3 + 4 + 5 = 4 + 3 + 5 B.-2(5 + 7) = -2(7 + 5) C.3(-8 + 1) = 3(-8) + 3(1) D.6[(7)(-2)] = [(6)(7)](-2)

OpenStudy (anonymous):

C

OpenStudy (anonymous):

because the distributive property needs the number outside to be multiplied to the numbers inside

OpenStudy (anonymous):

It's C. This is Distributive Property. \[4 \rightarrow \rightarrow\left( 3 \times 4 \right) = 4\times3 \times 4\times4 = 12\times 16. \]